# Mitigating Content Effects on Reasoning in Language Models through Fine-Grained Activation Steering

### v1

@@ 4 Empirical Evaluation
We evaluate the steering performance on three model families, covering different spans of model sizes: Llama (3.2-1b-it, 3.2-3b-it, 3.1-8b) grattafiori2024llama , Gemma-2 (2b-it, Gemma-2-9b-it) team2024gemma , Qwen 2.5 (1.5b-it, 3b-it, 7b-it) bai2023qwen . We use the instruction-tuned version of each model and evaluate both performance in zero-shot setting and in-context learning (ICL) via few-shot prompts, providing a total of 4 random examples from the training set.
Figure 2: Linear probing results for formal validity on Llama-3.1 8b. The probing experiments reveal that the information for both validity and plausibility is maximally localised in later layers, peaking at the last third quarter of the layers in the residual stream across different LLMs (see Appendix).
@@
To inform subsequent test-time interventions and steering, we performed a preliminary observational study through linear probing belinkov2022probing ; ferreira2021does to identify where information about the validity and plausibility of arguments might be encoded within the models. To this end, we employ a linear layer on top of the frozen activations of the models after processing a syllogistic argument, classifying whether the argument is valid/invalid and plausible/implausible. Figure 2 shows the results for formal validity on Llama 3 8B instruct (the results for plausibility and other models are reported in the Appendix). Overall, the probing experiments reveal that the information for both validity and plausibility is maximally localised in later layers, peaking at the last third quarter of the layers in the residual stream across different LLMs. Therefore, subsequent steering methods intervene on the last third quarter of layers at the last input token position.
@@
We adopt different evaluation metrics to compute the effect of steering on syllogistic reasoning. First, we compute the accuracy (ACC) of each model when assessing the validity of the syllogistic arguments in the test set. In addition, we measure the content effect (CE) based on the difference in accuracy on different subsets of the test set. In particular, we measure both the cross-plausibility CE as the difference in overall accuracy between plausible and implausible arguments, as well as the intra-plausibility CE as the difference in accuracy between valid and invalid arguments given a fixed plausibility value. The overall CE reported in the experiments is computed as the average of cross and intra-plausibility CE. Finally, we report the ACC/CE ratio, since the target objective is to obtain maximal accuracy on formal reasoning with minimal content effect.
@@
$\alpha=0$ $\ \alpha\neq 0$
Model Size Acc $\uparrow$ CE $\downarrow$ Acc/CE $\uparrow$ Acc $\uparrow$ CE $\downarrow$ Acc/CE $\uparrow$ $\alpha_{best}$ $\Delta\%$
Zero-shot
Llama 3.2 1b 58.17 44.04 1.32 73.56 6.35 11.58 -0.9 777.27
3b 77.79 17.50 4.45 77.79 17.50 4.45 0.0 0.00
Llama 3.1 8b 78.27 30.77 2.54 85.10 14.04 6.06 0.9 138.58
Gemma 2 2b 73.27 32.43 2.26 74.13 20.83 3.56 1.8 57.52
9b 85.00 8.46 10.05 83.27 1.92 43.37 0.6 331.54
Qwen 2.5 1.5b 75.67 14.42 5.25 77.79 12.88 6.04 0.3 15.05
3b 85.29 7.12 11.99 85.29 7.12 11.99 0.0 0.00
7b 88.85 5.39 16.48 89.90 0.96 93.65 -1.5 468.26
ICL
Llama 3.2 1b 57.21 45.70 1.25 66.44 6.08 4.94 -1.5 295.2
3b 72.79 28.14 2.59 78.27 22.31 3.51 0.3 35.52
Llama 3.1 8b 40.58 20.26 2.00 30.67 14.81 2.07 0.3 3.5
Gemma 2 2b 69.42 14.23 4.88 70.00 12.31 5.69 -0.3 16.6
9b 84.61 15.25 5.54 80.38 3.33 24.14 0.3 335.74
Qwen 2.5 1.5b 51.92 65.12 0.80 72.69 32.18 2.26 -1.2 182.5
3b 86.44 13.91 6.21 86.63 4.80 18.02 0.6 190.18
7b 89.80 9.36 9.60 89.90 4.81 18.70 0.9 94.79
Table 1: Results of contrastive steering with static values of $\alpha$. The table shows the results obtained with the best value of $\alpha$ based on accuracy and content effect ratio (ACC/CE) selected from the interval $[-3.0,3.0]$. A value of $\alpha_{best}=0$ indicates that the best results are achieved without steering. $\Delta\%$ represents the relative improvement of Acc/CE for each model. We found that contrastive steering is highly effective for most models, except Llama 3.2 3b and Qwen 2.5 3b.
4.1 Contrastive Activation Steering
We perform experiments on a test set of 2400 examples equally distributed across different types of arguments and schemes. We investigate the effect of steering varying the value of $\alpha$ between -3 and 3. The results are reported in Table 1. We summarise the main results and observations below.
Effectiveness of contrastive steering.
The results reveal that contrastive steering is effective for improving ACC/CE on most of the tested models in both zero-shot and ICL settings. Notably, contrastive steering has the highest impact on Llama 3.2 1b with a relative improvement of ACC/CE of up to 777.27%. A substantial improvement can be observed across different families and sizes of models (in particular for Gemma 2 9b and Qwen 2.5 7 via zero-shot). Moreover, we found that for most models, steering for content effect not only improves CE, but also contributes to significant improvements in accuracy on the syllogistic reasoning task (e.g., from 58.17% to 73.56% with Llama 1b). At the same time, steering with a static value of $\alpha$ seems to be ineffective on two zero-shot models – i.e., Llama 3.2 3b and Qwen 2.5 3b.
Steering outperforms ICL.
The scaling parameter $\alpha$ enables explicit steering control.
In order to investigate why steering is not effective on Llama 3.2 3b and Qwen 2.5 3b, we study the detailed dynamics emerging with different values of $\alpha$ (Figure 3, top). Here, we observe that, despite being ineffective on some models, contrastive steering can be used to explicitly control the accuracy achieved on valid and invalid arguments by varying the sign of $\alpha$. Specifically, the results show that setting $\alpha<0$ generally improves accuracy on valid arguments, while $\alpha>0$ improves accuracy on invalid arguments. This observation motivates us to explore conditional steering techniques to dynamically determine the value of $\alpha$ and attempt to steer unresponsive models.
(a) Acc – Llama 1b
(b) CE – Llama 1b
(c) Acc/CE – Llama 1b
(d) Acc – Llama 3b
(e) CE – Llama 3b
(f) Acc/CE – Llama 3b
(g) Acc – K-CAST
(h) CE – K-CAST
(i) Acc/CE – K-CAST
4.2 Conditional Activation Steering
Model Size Acc $\uparrow$ CE $\downarrow$ Acc/CE $\uparrow$ $\Delta\%$
No Steering
@@ 4.2 Conditional Activation Steering
Table 2: Results of conditional steering on models that are unresponsive to static contrastive steering – i.e, Llama 3.2 3b and Qwen 2.5 7b. We found that both CAST and K-CAST effectively improve ACC/CE for both models.
Computing Condition Vectors.
Motivated by the observation that the sign of $\alpha$ can enable explicit control on the task, we compute condition vectors to identify whether a given model is processing a valid or an invalid argument from the internal activations and then modulate the parameter $\alpha$ accordingly (i.e., setting $\alpha<0$ if the condition is valid, and $\alpha>0$ otherwise). To this end, we collect condition activation vectors for validity from the training set (separating valid from invalid arguments) and experiment with both CAST and K-CAST.
Conditional steering is effective on unresponsive models.
We found that both CAST and K-CAST are effective in improving ACC/CE for both Llama 3b and Qwen 3b (see Table 2). Moreover, while the results on Qwen show that CAST and K-CAST have a similar effect, the results on Llama reveal that K-CAST is significanlty more effective, leading to an absolute increase in accuracy of up to 15%. Figure 3 (bottom) shows the impact of K-CAST on Llama 3b with different values of $\alpha$ (with the sign dynamically determined at inference time).
4.3 Robustness to Prompt Perturbations
To test the robustness of steering performance to prompt perturbations mizrahi-etal-2024-state , we construct a set of prompt variants by employing instruction templates different from those used in the training set (i.e. instruction template paraphrasing). Following mizrahi-etal-2024-state , we employ two prompting strategies proven effective in prior research: (1) Instruction template rephrasing: we use GPT-4.5 to paraphrase a seed instruction template lester2021power ; gonen2022demystifying ; honovich-etal-2023-unnatural ; (2) Instruction induction: inspired by honovich-etal-2023-instruction , we provide five input-output pairs and ask GPT-4.5 to generate the possible instructions. Given a set of prompt variations (see Appendix), we compute the steering vectors using the original prompt and randomly select a variant at inference time.
Steering is robust to prompt variations.
The results in Figure 4 on Llama 1b (i.e., the model with the best ACC/CE improvement on the original prompt) reveal that, despite specific variation in the values of $\alpha_{best}$ and some noise deriving from prompt variation, the overall effectiveness of steering remains unaltered. A similar trend is also observable for other models (see Appendix).
@@ 4.3 Robustness to Prompt Perturbations
4.4 Impact of Steering on Non-Target Capabilities
@@
We draw 2,000 examples per language from the C4 dataset raffel2020exploring and compute the average perplexity on causal language modeling over sequences of length 1,024. Table 3 shows that content-effect steering leaves multilingual modeling nearly intact. For example, on English text, Llama 3.2 1b model’s perplexity changes only from 24.29 (baseline) to 24.77 (steered). Gemma exhibits the most significant relative increase, but even there, the gap remains small; across all languages and model sizes, perplexity deviations stay within a few percent.
@@
We test the steering impact on two reasoning tasks (i.e., ProntoQA PrOntoQA and Rulebreakers chan2024rulebreakers ) that were not presented during the steering modulating process. ProntoQA is a synthetic question-answering task designed to test deductive reasoning on multiple natural language premises. Rulebreakers is a question-answering task for evaluating LLMs ability to distinguish whether logical entailment diverges from factually acceptable inference. Table 3 show that adopting the steering vectors computed on syllogisms generalise well, especially on ProntoQA with Llama and Qwen (+8.1% and +5.2%), whereas Gemma experiences a substantial performance drop on both tasks, most notably a 16.1% decrease on ProntoQA (this aligns with the higher drops observed on language modeling). These findings underscore both the promise of steering for enhancing targeted reasoning and the persistent challenge of achieving robust OOD generalization.
@@ 5 Conclusion
This paper investigated the problem of mitigating content biases on formal reasoning in LLMs through activation steering. Specifically, we curated a controlled syllogistic reasoning dataset to disentangle formal validity from content plausibility and systematically tested both static and conditional steering techniques on different families of LLMs.
Our results indicate that contrastive steering is particularly effective in reducing content effect and improving accuracy in formal reasoning, with a newly introduced kNN-based conditional steering method being able to achieve up to 15% absolute accuracy improvement.
In general, this paper demonstrates that activation-level interventions offer a scalable inference-time strategy for enhancing the robustness of LLMs, contributing towards more systematic and unbiased formal reasoning. In future work, we plan to explore how to enable better OOD generalization via steering for improving and debiasing more general reasoning capabilities in LLMs in different tasks and domains. To this end, we believe that conditional steering represents a promising solution via the dynamic adaptation of the steering modalities at test time. For this, different conditions could be explored to dynamically target different forms of reasoning.
6 Limitations

### v2

@@ 4 Empirical Evaluation
We evaluate the steering performance on three model families, covering different spans of model sizes: Llama (3.2-1b-it, 3.2-3b-it, 3.1-8b) (Grattafiori et al. 2024), Gemma-2 (2b-it, Gemma-2-9b-it) (Team et al. 2024), Qwen 2.5 (1.5b-it, 3b-it, 7b-it) (Bai et al. 2023). We use the instruction-tuned version of each model and evaluate both performance in zero-shot setting and in-context learning (ICL) via few-shot prompts, providing a total of 4 random examples from the training set.
@@
To inform subsequent test-time interventions and steering, we performed a preliminary observational study through linear probing (Belinkov 2022; Ferreira et al. 2021) to identify where information about the validity and plausibility of arguments might be encoded within the models. To this end, we employ a linear layer on top of the models’ frozen activations after processing a syllogistic argument, classifying whether the argument is valid/invalid and plausible/implausible. Overall, the probing experiments reveal that the information for validity and plausibility is maximally localised in later layers in the residual stream, consistently peaking at the third quarter of the layers across different LLMs (the detailed results for probing formal validity and plausibility can be found in the supplementary material). Therefore, following the insights from probing, subsequent steering methods intervene within the third quarter of layers at the residual stream corresponding to the last input token position.
@@
We adopt different evaluation metrics to compute the effect of steering on syllogistic reasoning. First, we compute the accuracy (ACC) of each model when assessing the validity of the syllogistic arguments in the test set. In addition, we measure the content effect (CE) based on the difference in accuracy on different subsets of the test set. In particular, we measure both the cross-plausibility CE as the difference in overall accuracy between plausible and implausible arguments, as well as the intra-plausibility CE as the difference in accuracy between valid and invalid arguments given a fixed plausibility value. The overall CE reported in the experiments is computed as the average of cross and intra-plausibility CE. Finally, we report the Acc/CE ratio, as the objective is to obtain maximal accuracy on formal reasoning with minimal content effect.
@@
Model Size Acc $\uparrow$ CE $\downarrow$ Acc/CE $\uparrow$ $\Delta_{rel}^{Acc/CE}$
@@
Table 2: Results of conditional steering on models that are unresponsive to static contrastive steering – i.e, Llama 3.2 3b and Qwen 2.5 7b. We found that both CAST and K-CAST effectively improve Acc/CE.
@@
4.1 Contrastive Activation Steering
We perform experiments on a test set of 2400 examples equally distributed across different syllogistic schemes. We investigate the effect of steering by varying the value of $\alpha$ between -3 and 3. The results are reported in Table 1.
Effectiveness of contrastive steering.
The results reveal that contrastive steering is effective for improving Acc/CE on most of the tested models in both zero-shot and ICL settings. Notably, contrastive steering has the highest impact on Llama 3.2 1b with a relative improvement of Acc/CE of up to 777.27%. A substantial improvement can be observed across different families and sizes of models (in particular for Gemma 2 9b and Qwen 2.5 7 via zero-shot). Moreover, we found that for most models, steering for content effect not only improves CE, but also contributes to significant improvements in accuracy on the syllogistic reasoning task (e.g., from 58.17% to 73.56% with Llama 1b). At the same time, steering with a static value of $\alpha$ is ineffective on two zero-shot models – i.e., Llama 3.2 3b and Qwen 2.5 3b.
Steering outperforms ICL.
The scaling parameter $\alpha$ enables explicit steering control.
In order to investigate why steering is not effective on Llama 3.2 3b and Qwen 2.5 3b, we study the detailed dynamics emerging with different values of $\alpha$. Here, we observe that, despite being ineffective on some models, contrastive steering can be used to explicitly control the accuracy achieved on valid and invalid arguments by varying the sign of $\alpha$. Specifically, the results show that setting $\alpha<0$ generally improves accuracy on valid arguments, while $\alpha>0$ improves accuracy on invalid arguments. This observation motivates us to explore conditional steering techniques to dynamically determine the value of $\alpha$ and attempt to steer unresponsive models.
4.2 Conditional Activation Steering
Computing Condition Vectors.
Motivated by the observation that the sign of $\alpha$ enables explicit control, we compute condition vectors to identify whether a given model is processing a valid or an invalid argument from the internal activations and then modulate the parameter $\alpha$ accordingly (i.e., setting $\alpha<0$ if valid, and $\alpha>0$ otherwise). To this end, we collect condition activation vectors for validity from the training set and experiment with both CAST and K-CAST.
Conditional steering is effective on unresponsive models.
We found that both CAST and K-CAST are effective in improving Acc/CE for both Llama 3b and Qwen 3b (see Table 2). Moreover, while the results on Qwen show that CAST and K-CAST have a similar effect, the results on Llama reveal that K-CAST is significantly more effective, leading to an absolute increase in accuracy of up to 15%. Figure 3 (b) shows the impact of K-CAST on Llama 3b with different values of $\alpha$ (with the sign dynamically determined).
Figure 4: Robustness of steering to prompt variations on Llama 1b (i.e. Acc/CE). The results reveal that, despite specific variations deriving from perturbations applied at test time, the overall effectiveness of steering remains unaltered.
4.3 Robustness to Prompt Perturbations
To test the robustness of steering performance to prompt perturbations (Mizrahi et al. 2024), we construct a set of prompt variants by employing instruction templates different from those used in the training set (i.e. instruction template paraphrasing). Following (Mizrahi et al. 2024), we employ two prompting strategies proven effective in prior research: (1) Instruction template rephrasing: we use GPT-4.5 to paraphrase a seed instruction template (Lester et al. 2021; Gonen et al. 2022; Honovich et al. 2023a); (2) Instruction induction: inspired by (Honovich et al. 2023b), we provide five input-output pairs and ask GPT-4.5 to generate the possible instructions. Given a set of prompt variations (see supplementary material), we compute the steering vectors using the original prompt and randomly select a variant at inference time.
Steering is robust to prompt variations.
The results in Figure 9 on Llama 1b (i.e., the model with the best Acc/CE improvement on the original prompt) reveal that, despite specific variation in the values of $\alpha_{best}$ and some noise deriving from prompt variation, the overall effectiveness of steering remains unaltered. A similar trend is also observable for other models (see supplementary material).
4.4 Impact on Non-Target Capabilities
@@
We draw 2,000 examples per language from the C4 dataset (Raffel et al. 2020) and compute the average perplexity on causal language modeling over sequences of length 1,024. Table 3 shows that content-effect steering leaves multilingual modeling nearly intact. For example, on English text, Llama 3.2 1b model’s perplexity changes only from 24.29 (baseline) to 24.77 (steered). Gemma exhibits the most significant relative increase, but even there, the gap remains small; across all languages and model sizes, perplexity deviations stay within a few percent.
@@
We test the steering impact on two reasoning tasks (i.e., ProntoQA (Saparov and He 2023) and Rulebreakers (Chan et al. 2024)) that were not presented during the steering modulating process. ProntoQA is a synthetic task designed to test deductive reasoning on multiple natural language premises. Rulebreakers is a task for evaluating LLMs ability to distinguish whether logical entailment diverges from factually acceptable inference. Table 3 show that adopting the steering vectors computed on syllogisms generalise well, especially on ProntoQA with Llama and Qwen (+8.1% and +5.2%), whereas Gemma experiences a substantial performance drop on both tasks, most notably a 16.1% decrease on ProntoQA (this aligns with the higher drops observed on language modeling).
These findings underscore both the promise of steering for enhancing targeted reasoning and the persistent challenge of achieving a complete and robust OOD generalization.
@@ 5 Conclusion
This paper investigated how to mitigate content effects in LLMs’ reasoning. We systematically tested static and conditional activation steering techniques to disentangle formal validity from content plausibility in syllogistic reasoning. Our results indicate that steering is particularly effective in reducing content effect and improving accuracy in formal reasoning. In general, this paper demonstrates that activation-level interventions offer a scalable inference-time strategy and can contribute towards more systematic and unbiased formal reasoning in LLMs.
Acknowledgments
This work was partially funded by the SNSF project NeuMath (200021_204617), by the CRUK National Biomarker Centre, and supported by the Manchester Experimental Cancer Medicine Centre and the NIHR Manchester Biomedical Research Centre.
Appendix A Experimental Setup
We run all the steering experiments on a single A100 GPU. For probing, the graphs report the variance obtained with different random seeds (across 10 runs). For steering, We set the seed to 0 to enable a reproducible behaviour.
Appendix B Syllogistic Schemes
Here is the list of syllogistic schemes adopted to construct the dataset. The schemes are aligned with the ones used in previous work on syllogistic reasoning with LLMs (Bertolazzi et al. 2024a).
Schema: AA1
Premise 1: All <A> are <B>
Premise 2: All <B> are <C>
Conclusions: All <A> are <C> | some <A> are <C> | some <C> are <A>
Schema: AA2
Premise 1: All <B> are <A>
Premise 2: All <C> are <B>
Conclusions: All <C> are <A> | some <A> are <C> | some <C> are <A>
Schema: AA4
Premise 1: All <B> are <A>
Premise 2: All <B> are <C>
Conclusions: Some <A> are <C> | some <C> are <A>
Schema: AI2
Premise 1: All <B> are <A>
Premise 2: Some <C> are <B>
Conclusions: Some <A> are <C> | some <C> are <A>
Schema: AI4
Premise 1: All <B> are <A>
Premise 2: Some <B> are <C>
Conclusions: Some <A> are <C> | some <C> are <A>
Schema: AO3
Premise 1: All <A> are <B>
Premise 2: Some <C> are not <B>
Conclusions: Some <C> are not <A>
Schema: AO4
Premise 1: All <B> are <A>
Premise 2: Some <B> are not <C>
Conclusions: Some <A> are not <C>
Schema: AE1
Premise 1: All <A> are <B>
Premise 2: No <B> are <C>
Conclusions: No <A> are <C> | no <C> are <A> | some <A> are not <C> | some <C> are not <A>
Schema: AE2
Premise 1: All <B> are <A>
Premise 2: No <C> are <B>
Conclusions: Some <A> are not <C>
Schema: AE3
Premise 1: All <A> are <B>
Premise 2: No <C> are <B>
Conclusions: No <A> are <C> | no <C> are <A> | some <A> are not <C> | some <C> are not <A>
Schema: AE4
Premise 1: All <B> are <A>
Premise 2: No <B> are <C>
Conclusions: Some <A> are not <C>
Schema: IA1
Premise 1: Some <A> are <B>
Premise 2: All <B> are <C>
Conclusions: Some <A> are <C> | some <C> are <A>
Schema: IA4
Premise 1: Some <B> are <A>
Premise 2: All <B> are <C>
Conclusions: Some <A> are <C> | some <C> are <A>
Schema: IE1
Premise 1: Some <A> are <B>
Premise 2: No <B> are <C>
Conclusions: Some <A> are not <C>
Schema: IE2
Premise 1: Some <B> are <A>
Premise 2: No <C> are <B>
Conclusions: Some <A> are not <C>
Schema: IE3
Premise 1: Some <A> are <B>
Premise 2: No <C> are <B>
Conclusions: Some <A> are not <C>
Schema: IE4
Premise 1: Some <B> are <A>
Premise 2: No <B> are <C>
Conclusions: Some <A> are not <C>
Schema: OA3
Premise 1: Some <A> are not <B>
Premise 2: All <C> are <B>
Conclusions: Some <A> are not <C>
Schema: OA4
Premise 1: Some <B> are not <A>
Premise 2: All <B> are <C>
Conclusions: Some <C> are not <A>
Schema: EA1
Premise 1: No <A> are <B>
Premise 2: All <B> are <C>
Conclusions: Some <C> are not <A>
Schema: EA2
Premise 1: No <B> are <A>
Premise 2: All <C> are <B>
Conclusions: No <A> are <C> | no <C> are <A> | some <A> are not <C> | some <C> are not <A>
Schema: EA3
Premise 1: No <A> are <B>
Premise 2: All <C> are <B>
Conclusions: No <A> are <C> | no <C> are <A> |some <A> are not <C> | some <C> are not <A>
Schema: EA4
Premise 1: No <B> are <A>
Premise 2: All <B> are <C>
Conclusions: Some <C> are not <A>
Schema: EI1
Premise 1: No <A> are <B>
Premise 2: Some <B> are <C>
Conclusions: Some <C> are not <A>
Schema: EI2
Premise 1: No <B> are <A>
Premise 2: Some <C> are <B>
Conclusions: Some <C> are not <A>
Schema: EI3
Premise 1: No <A> are <B>
Premise 2: Some <C> are <B>
Conclusions: Some <C> are not <A>
Schema: EI4
Premise 1: No <B> are <A>
Premise 2: Some <B> are <C>
Conclusions: Some <C> are not <A>
Appendix C Prompts Templates
Variation 0:
Given the premises, evaluate
the validity of the conclusion.
Premises:
entry[‘premise1’].
entry[‘premise2’].
Conclusion: entry[‘conclusion’].
The conclusion is
Variation 1:
Carefully evaluate the validity
of the following logical argument and
answerthe argument is logically valid
or the argument is logically invalid’.
Premise1: entry[‘premise1’].
Premise2: entry[‘premise2’].
Conclusion: entry[‘conclusion’].
The argument is logically
Variation 2:
Analyze the formal logical structure of
the argument below, then indicate
whether it is valid or invalid.
[…]
The logical structure is
Variation 3:
Assess carefully whether the argument
presented below is logically valid or
invalid based on the given premises and
conclusion.
1. entry[‘premise1’]
2. entry[‘premise2’]
Conclusion:
entry[‘conclusion’]
The argument is logically
Variation 4:
Examine the following argument,
generating "valid" if the conclusion
logically follows from the premises
provided, "invalid" otherwise.
- Premise 1: entry[‘premise1’]
- Premise 2: entry[‘premise2’]
Conclusion: entry[‘conclusion’]
The argument is
Variation 5:
Given the two premises and the
conclusion below, judge carefully
whether the logical argument is valid or
invalid.
Premise 1: entry[‘premise1’]
Premise 2: entry[‘premise2’]
Conclusion: entry[‘conclusion’]
The logical argument is
Variation 6:
Evaluate the following logical argument
carefully and decide whether the
provided conclusion is formally valid or
invalid given the two premises.
Premise 1: entry[‘premise1’]
Premise 2: entry[‘premise2’]
Conclusion: entry[‘conclusion’]
The conclusion is formally
Appendix D Linear Probing Results
Results are reported in Figure 2 and Figure 5.
Appendix E Static and Conditional Steering
Detailed results of static and conditional steering on Llama are reported in Figure 6. A comparison between CAST and K-CAST is reported in Figure 7 and Figure 8.
Appendix F Robustness to Prompt Perturbations
Detailed results are reported in Figure 9 and Figure 10.
Appendix G Limitations